OpenStudy (anonymous):

Solve the system of linear equations by elimination. check your solution. x+y=3 and x-y=1

OpenStudy (anonymous):

You have +y in the first equation, And -y in the second I think a good way to eliminate y would be to add the two equations together x + x + y - y = 3 + 1 2 x = 4 x = ?